# A Survey on Text-Driven 360-Degree Panorama Generation

### v1

@@ 4.1 Text-Only Generation
Table 2: Summary of Text-Only Generation. ‘LDM-B’ indicates whether the method is based on latent diffusion models. ‘TF’ specifies if it is training-free. ‘Code’ denotes whether both the source code and the pre-trained model checkpoint are publicly accessible.
4.2 Text-Driven NFoV Outpainting
This paradigm enhances user control by conditioning the generation process on both textual prompts and initial narrow NFoV images. The NFoV image, representing a limited portion of the scene, provides a visual starting point, which the generative model then extends into a complete panorama guided by the textual description. Tab. 3 offers a summary of representative text-driven NFoV outpainting approaches.
An early attempt in this paradigm is ImmerseGAN Dastjerdi et al. (2022), which uses a GAN-based inpainting architecture for this task. To achieve text-guided outpainting, ImmerseGAN adopts a pre-trained discriminative network to produce a latent vector representing the given textual description. This latent vector subsequently guides the generator in producing a 360-degree panorama that is semantically consistent with the text prompt. More recent approaches have primarily focused on leveraging the power of pre-trained latent diffusion models (LDMs) for their strong image generation priors acquired from training on large-scale datasets. These methods can be broadly classified into autoregressive-based (AR-based) and non-autoregressive-based (NAR-based) approaches according to their underlying frameworks.
AR-Based.
AOG-Net Lu et al. (2024) introduces an autoregressive framework, building upon a pre-trained LDM Rombach et al. (2022), to progressively outpaint NFoV images into complete panoramas under textual guidance. This stepwise approach enhances the generation of fine-grained visual content and improves alignment with the input textual descriptions. Moreover, AOG-Net incorporates a global-local conditioning mechanism to integrate text, omni-visual, NFoV and omni-geometry guidances at each autoregression step, thereby ensuring consistency across the generated panorama. AOG-Net is trained on Laval Indoor Gardner et al. (2017) and Laval Outdoor Hold-Geoffroy et al. (2019) for indoor and outdoor scenarios, respectively. Following the training dataset settings in AOG-Net, OPa-Ma Gao et al. (2024) uses a pre-trained LDM with Mamba Gu and Dao (2024), a state-space model known for its efficiency in handling long sequences, to iteratively outpaint local regions in each step. It introduces two modules: the Visual-textual Consistency Refiner, which enhances input utilization during generation, and the Global-local Mamba Adapter, which ensures global coherence across the generated panorama.
NAR-Based.
PanoDiff Wang et al. (2023), the first LDM-based method for text-driven NFoV outpainting, is trained on SUN360 Xiao et al. (2012). It first converts the input NFoV images into partial panoramas with visibility masks, and then employs a ControlNet-based LDM Zhang et al. (2023) for text-guided panorama completion. To ensure geometric continuity at the borders of the generated panorama, PanoDiff further implements a circular padding scheme during inference. Similarly, Diffusion360 Feng et al. (2023) adopts a ControlNet-based LDM Zhang et al. (2023) to generate 360-degree panoramas from perspective images and textual descriptions. However, instead of circular padding, Diffusion360 leverages a circular blending strategy during the denoising and VAE decoding stages for improved boundary continuity of the generated 360-degree panorama.
Unlike previous methods Wang et al. (2023); Feng et al. (2023) that produce 360-degree panoramas using equirectangular representation, CubeDiff Kalischek et al. (2025), inspired by multi-view diffusion models Shi et al. (2024); Tang et al. (2023), generates 360-degree panoramas in cubemap format. This cubemap representation enables CubeDiff to more effectively leverage the diffusion priors learned by the LDM from extensive perspective images during the generation process. CubeDiff fine-tunes a pre-trained LDM on a mixed dataset of Structured3D Zheng et al. (2020), Pano360 Kocabas et al. (2021), Polyhaven Poly Haven (2025), and Humus Persson (2025), using a single conditional view (NFoV image) and textual embeddings as input. To model inter-face relationships within the cubemap, CubeDiff inflates all 2D attention layers in the LDM, enabling effective learning of inter-view dependencies.
Method Publication LDM-B TF Code
ImmerseGAN 3DV 2022 $\times$ $\times$ $\times$
PanoDiff MM 2023 $\checkmark$ $\times$ $\checkmark$
Diffusion360 arxiv 2023 $\checkmark$ $\times$ $\checkmark$
AOG-Net AAAI 2024 $\checkmark$ $\times$ $\times$
OPa-Ma arxiv 2024 $\checkmark$ $\times$ $\times$
CubeDiff ICLR 2025 $\checkmark$ $\times$ $\times$
@@ 4.2 Text-Driven NFoV Outpainting
Table 3: Summary of Text-Driven NFoV Outpainting. For explanations of the ‘LDM-B’, ‘TF’ and ‘Code’ columns, see Tab. 2.
Method FID $\downarrow$ KID (×$\times 10^{-2}$) $\downarrow$ IS $\uparrow$ CS $\uparrow$ FAED $\downarrow$ OmniFID $\downarrow$ DS $\downarrow$ Inference (s) $\downarrow$
Text2Light 72.63 1.54 5.35 19.20 18.10 99.81 5.38 33
Diffusion360 70.32 2.00 5.29 18.74 12.43 92.23 0.94 3
StitchDiffusion 76.69 2.04 7.36 19.20 15.58 108.63 1.07 28
PanFusion 61.23 1.07 6.16 18.96 13.16 92.22 0.85 30
PanoDiff 65.94 2.44 4.72 19.02 10.24 122.30 1.10 48
Diffusion360 64.19 2.05 4.53 17.92 5.50 101.39 0.72 4
Table 4: Quantitative Comparison of Representative Text-Driven 360-Degree Panorama Generation. The first block of rows are for methods in the paradigm of Text-Only Generation, while the second block of rows are for Text-Driven NFoV Outpainting. We use metrics outlined in Sec. 3.3 for comprehensive evaluation. The inference time, required by each method to generate a 1024$\times$512 360-degree panorama, is also reported. The best and second-best results are highlighted for the two paradigms, respectively.
4.3 Quantitative Comparison
@@
To systematically evaluate strengths and weaknesses of representative methods of different kinds, we benchmark Text-Only Generation (Sec. 4.1) and Text-Driven NFoV Outpainting (Sec. 4.2) methods with publicly available inference codes and model checkpoints. For Text-Only Generation, we compare Text2Light Chen et al. (2022), Diffusion360 Feng et al. (2023), StitchDiffusion Wang et al. (2024) and PanFusion Zhang et al. (2024). For Text-Driven NFoV Outpainting, we compare PanoDiff Wang et al. (2023) and Diffusion360. We conduct a comprehensive comparison using metrics outlined in Sec. 3.3.

### v2

@@ III-C Evaluation Metrics
TABLE II: Summary of Text-Only Generation. ‘LDM-B’ indicates whether the method is based on latent diffusion models. ‘TF’ specifies if it is training-free. ‘Code’ denotes whether both the source code and the pre-trained model checkpoint are publicly accessible. ‘N/A’ means not applicable.
Method Publication LDM-B TF Code Training Datasets Representation
Text2Light [56] TOG 2022 $\times$ $\times$ $\checkmark$ Polyhaven [5], Laval Indoor [45], Laval Outdoor [46] ERP
iHDRI [70], HDRI Skies [68], HDRMaps [69]
Diffusion360 [58] arxiv 2023 $\checkmark$ $\times$ $\checkmark$ SUN360 [48] ERP
StitchDiffusion [57] WACV 2024 $\checkmark$ $\times$ $\checkmark$ Polyhaven [5] ERP
PanFusion [59] CVPR 2024 $\checkmark$ $\times$ $\checkmark$ Matterport3D [49] ERP
PanoFree [41] ECCV 2024 $\checkmark$ $\checkmark$ $\times$ N/A MPP
DiffPano [40] NeurIPS 2024 $\checkmark$ $\times$ $\times$ Habitat Matterport 3D [15] ERP
$\text{Omni}^{2}$ [23] arxiv 2025 $\times$ $\times$ $\times$ SUN360 [48], Structured3D [47] MPP
SphereDiff [22] arxiv 2025 $\checkmark$ $\checkmark$ $\times$ N/A Spherical
PAR [25] arxiv 2025 $\times$ $\times$ $\times$ Matterport3D [49] ERP
UniPano [80] arxiv 2025 $\checkmark$ $\times$ $\times$ Matterport3D [49] ERP
The universal metrics provide general assessments but often fail to capture nuances and geometric distortions unique to 360-degree panoramas. Panorama-specific metrics designed specifically for ERP-based 360-degree panoramic images include Fréchet Auto-Encoder Distance (FAED), Omnidirectional FID (OmniFID), Discontinuity Score (DS), and Distortion-perception FID (Distort-FID).
FAED [50] computes Fréchet distances between features extracted from generated and real panoramas. Unlike FID, it employs an autoencoder [34] specifically trained on 360-degree panoramic images. Lower FAED scores reflect better perceptual and geometric quality tailored to the unique panoramic properties.
OmniFID [51] adapts FID to specifically evaluate 360-degree panoramas. It converts equirectangular panoramas into cubemap representations and calculates FID across three disjoint subsets of cubemap faces, averaging the results. Lower OmniFID scores indicate higher geometric fidelity in 360-degree panorama generation.
DS [51] measures the seam alignment across the borders of generated panoramas by applying a kernel-based edge detection algorithm. Lower discontinuity scores correspond to fewer visible seam artifacts, indicating better perceived consistencies across the seam.
Distort-FID [24] measures the distance between feature distributions of generated and real images based on a distortion-perception CLIP [24] that is fine-tuned on 360-degree panoramic images. Lower Distort-FID scores reflect better distortion accuracy of the generated 360-degree panoramic images.
We provide a comprehensive comparison of state-of-the-art methods, introduced in the following section, using the outlined metrics in Sec. IV-C2.
IV State-of-the-Art Methods
Existing text-driven 360-degree panorama generation methods can be broadly categorized into two paradigms according to input modalities: (a) Text-Only Generation aims to synthesize 360-degree panoramas from textual prompts only, while (b) Text-Driven Narrow Field-of-View (NFoV) Outpainting leverages both textual descriptions and initial NFoV images to guide the generation process, offering enhanced user control. Fig. 5 provides an intuitive illustration for both paradigms. We detail the literature for both as follows.
IV-A Text-Only Generation
This paradigm focuses on synthesizing 360-degree panoramas from textual descriptions only. Tab. II provides a comparative overview of representative text-only methods. These methods can be broadly divided into training-based and training-free approaches.
IV-A1 Training-Based
Text2Light [56], an early notable effort, explored a hierarchical text-driven framework utilizing VQGAN [43] and CLIP [52] to address this challenge based on training data aggregated from multiple sources, such as Polyhaven [5], Laval Indoor [45] and Laval Outdoor [46]. Recently, the advent of latent diffusion models (LDMs) [63] for text-to-image synthesis marked a significant advancement, enabling more sophisticated 360-degree panorama generation techniques. LDMs are typically trained on vast datasets consisting of standard perspective images with a limited field of view and corresponding textual descriptions. Despite demonstrating robust capabilities in generating perspective images from text prompts, these models face significant difficulties when creating 360-degree panoramas with a complete 360°$\degree$$\times$180°$\degree$ field of view, which differ substantially from traditional perspective images.
To adapt pre-trained LDMs for 360-degree panorama synthesis, a common strategy is to fine-tune these models with specialized 360-degree panorama datasets. Diffusion360 [58] exemplifies this approach by leveraging the DreamBooth technique [38] to fine-tune a pre-trained LDM [63] on SUN360 [48]. To ensure geometric consistency of boundaries, Diffusion360 utilizes a circular blending strategy during both the denoising process and the VAE [74] decoding stage, effectively reducing seam artifacts. In addition, it introduces a super-resolution module to enable the generation of high-resolution (6114$\times$3072) 360-degree panoramas. Notably, the DreamBooth fine-tuning approach is computationally demanding and time-consuming because it generally fine-tunes a significant portion of the model parameters.
In contrast, LoRA [39] has recently gained attention as a parameter-efficient fine-tuning method. LoRA works by injecting trainable low-rank matrices into the pre-trained model’s weights, allowing for rapid adaptation to new tasks with minimal additional parameters. For example, StitchDiffusion [57] employs LoRA to fine-tune a pre-trained LDM on a dataset of paired 360-degree images (sourced from Polyhaven [5]) and corresponding textual descriptions generated using BLIP [17]. It further introduces a tailored generation method based on MultiDiffusion [55] to ensure boundary continuity in the generated panoramas.
@@ IV-A1 Training-Based
TABLE III: Summary of Text-Driven NFoV Outpainting. For explanations of the ‘LDM-B’, ‘TF’ and ‘Code’ columns, see Tab. II.
Method Publication LDM-B TF Code Training Datasets Representation
ImmerseGAN [14] 3DV 2022 $\times$ $\times$ $\times$ 360Cities [71] ERP
PanoDiff [42] MM 2023 $\checkmark$ $\times$ $\checkmark$ SUN360 [48] ERP
Diffusion360 [58] arxiv 2023 $\checkmark$ $\times$ $\checkmark$ SUN360 [48] ERP
AOG-Net [60] AAAI 2024 $\checkmark$ $\times$ $\times$ Laval Indoor [45], Laval Outdoor [46] ERP
OPa-Ma [4] arxiv 2024 $\checkmark$ $\times$ $\times$ Laval Indoor [45], Laval Outdoor [46] ERP
CubeDiff [44] ICLR 2025 $\checkmark$ $\times$ $\times$ Polyhaven [5], Humus [6], Structured3D [47] CMP
Pano360 [19]
PanoDecouple [24] CVPR 2025 $\checkmark$ $\times$ $\times$ SUN360 [48] ERP
$\text{Omni}^{2}$ [23] arxiv 2025 $\times$ $\times$ $\times$ SUN360 [48], Structured3D [47] MPP
PAR [25] arxiv 2025 $\times$ $\times$ $\times$ Matterport3D [49] ERP
Other works have similarly adopted the LoRA fine-tuning technique. PanFusion [59] devises a dual-branch diffusion model, trained on the Matterport3D [49] dataset, with separate LoRA layers to integrate both global panoramic and local perspective views. It introduces an equirectangular-perspective projection attention module to facilitate information exchange between the two branches, aiming to alleviate visual inconsistencies in the generated 360-degree panoramas. However, PanFusion’s output often exhibits blurriness at the top and bottom regions, due to its training dataset. To avoid this issue, DiffPano [40] utilizes the Habitat Matterport 3D dataset [15] to produce multi-view consistent 360-degree panoramas with clearer top and bottom details. For generating more precise textual descriptions, DiffPano adopts BLIP2 [18] and Llama2 [16] sequentially, resulting in a panoramic video-text dataset. Based on this dataset, it fine-tunes a pre-trained LDM [63] using LoRA for single-view text-driven 360-degree panorama generation. Furthermore, to enable multi-view 360-degree panorama generation based on text prompts and camera viewpoints, DiffPano proposes a spherical epipolar-aware attention module, which effectively extends single-view panorama generation to multi-view consistent panorama generation. Based on the observation that value and output weight matrices are crucial during the LoRA-based fine-tuning of cross-attention blocks within pre-trained LDMs for 360-degree panoramic image generation, UniPano [80] proposes a uni-branch framework for panorama generation. This framework exclusively fine-tunes these specific matrices using LoRA, while keeping the original query and key weight matrices in the cross-attention blocks frozen.
More recently, diverging from fine-tuning the commonly used pre-trained LDMs [63] (as in PanFusion [59] and DiffPano [40]), several studies have explored alternative pre-trained models for synthesizing 360-degree panoramic images from textual descriptions. Inspired by OmniGen [76], $\text{Omni}^{2}$ [23] adopted a diffusion model consisting of a VAE [74] and a pre-trained Transformer [75]. To adapt the pre-trained Transformer for synthesizing 360-degree panoramic images, LoRA fine-tuning technique is employed. Instead of processing the entire 360-degree panoramic image at once, $\text{Omni}^{2}$ generates six overlapping viewports. In the inference phase, these synthesized perspective images are integrated to reconstruct 360-degree panoramic images. Addressing the challenge that spatial distortions in ERP 360-degree panoramic images violate the identically and independently distributed (i.i.d.) Gaussian noise assumption inherent in many diffusion models, PAR [25], inspired by masked autoregressive modeling (MAR) [77], proposes an autoregressive modeling approach for text-based 360-degree panoramic image generation. This method is not constrained by the i.i.d. assumption. Specifically, PAR fine-tunes a pre-trained autoregressive model [78] on the Matterport3D dataset [49] and develops a dual-space circular padding technique to mitigate boundary discontinuities.
IV-A2 Training-Free
In contrast to training-based approaches, PanoFree [41], introduces a tuning-free multi-view image generation framework based on a pre-trained LDM [63]. Guided by textual descriptions, PanoFree leverages iterative warping and inpainting steps to produce multi-view perspective images, which are subsequently stitched into 360-degree panoramas, thus avoiding the need for specialized 360-degree panorama datasets. Unlike PanoFree, which operates on perspective latent representations, the recent SphereDiff [22] proposes leveraging a spherical latent representation for training-free generation of 360-degree panoramic images. Specifically, SphereDiff extends the MultiDiffusion [55] framework to this spherical latent space. To mitigate minor distortions arising from the spherical-to-perspective projection during its process, SphereDiff further incorporates a distortion-aware weighted averaging method.
IV-B Text-Driven NFoV Outpainting
This paradigm enhances user control by conditioning the 360-degree panorama generation process on both textual prompts and initial NFoV images. The NFoV image, representing a limited portion of the scene, serves as a visual starting point, which the generative model subsequently expands into a complete 360-degree panoramic image guided by the textual description. Tab. III offers a summary of representative text-driven NFoV outpainting approaches. These approaches can be broadly classified into non-autoregressive-based (NAR-based) and autoregressive-based (AR-based) methods according to their underlying frameworks.
IV-B1 NAR-Based
An early attempt in this paradigm, ImmerseGAN [14], utilized a GAN-based inpainting architecture [79] for this task. To achieve text-guided outpainting, ImmerseGAN adopted a pre-trained discriminative network to produce a latent vector representing the given textual description. This latent vector subsequently guided the generator to produce a 360-degree panorama semantically consistent with the text prompt. More recent approaches have primarily focused on leveraging the power of pre-trained latent diffusion models (LDMs) for their strong image generation priors acquired from training on large-scale datasets.
PanoDiff [42], the first LDM-based method for text-driven NFoV outpainting, was trained on the SUN360 [48] dataset. It initially converts the input NFoV images into partial panoramas with visibility masks, and then employs a ControlNet-based LDM [37] for text-guided panorama completion. To ensure geometric continuity at the borders of the generated panorama, PanoDiff further implements a circular padding scheme during inference. Similarly, Diffusion360 [58] adopts a ControlNet-based LDM [37] to generate 360-degree panoramas from perspective images and textual descriptions. However, instead of circular padding, Diffusion360 leverages a circular blending strategy during the denoising and VAE decoding stages for improved boundary continuity of the generated 360-degree panorama. Recognizing that a single network (as employed in PanoDiff and Diffusion360) often struggles to simultaneously learn the inherent 360-degree panoramic distortion and perform content completion, PanoDecouple [24] designs a decoupled diffusion model. This framework decouples the NFoV outpainting process into distortion guidance and content completion. Furthermore, it develops a distortion correction loss, leveraging the proposed Distort-CLIP, to explicitly constrain the distinct 360-degree panoramic distortion.
Depart from the aforementioned methods [24, 42, 58], which predominantly process and generate 360-degree panoramas using an equirectangular representation throughout their networks, several recent studies have explored leveraging alternative representations for 360-degree panoramic synthesis. CubeDiff [44], inspired by multi-view diffusion models [62, 61], generates 360-degree panoramas in cubemap format. This cubemap representation enables CubeDiff to more effectively leverage the diffusion priors learned by the LDM from extensive perspective images during the generation process. CubeDiff fine-tunes a pre-trained LDM on a mixed dataset of Structured3D [47], Pano360 [19], Polyhaven [5], and Humus [6], using a single conditional view (NFoV image) and textual embeddings as input. To model inter-face relationships within the cubemap, CubeDiff inflates all 2D attention layers in the LDM, enabling effective learning of inter-view dependencies. Similarly, $\text{Omni}^{2}$ [23] produces six overlapping perspective images, each with a 110°$\degree$$\times$110°$\degree$ field of view, using a Transformer [75] architecture fine-tuned with LoRA on the SUN360 [48] and Structured3D [47] datasets. These overlapping images are subsequently stitched together to synthesize the final 360-degree panoramic image.
Figure 6: Visual comparison of text-only generation methods. To facilitate the assessment of visual and semantic continuities at the image borders, we generate ”swapped” versions by exchanging the left and right halves of each generated image, and then use red dashed boxes in these swapped images to highlight the boundary regions.
IV-B2 AR-Based
AOG-Net [60] introduces an autoregressive framework, building upon a pre-trained LDM [63], to progressively outpaint NFoV images into complete panoramas under textual guidance. This stepwise approach enhances the generation of fine-grained visual content and improves alignment with the input textual descriptions. Moreover, AOG-Net incorporates a global-local conditioning mechanism to integrate text, omni-visual, NFoV and omni-geometry guidances at each autoregression step, thereby ensuring consistency across the generated panorama. AOG-Net is trained on Laval Indoor [45] and Laval Outdoor [46] for indoor and outdoor scenarios, respectively. Following the training dataset settings in AOG-Net, OPa-Ma [4] uses a pre-trained LDM with Mamba [3], a state-space model known for its efficiency in handling long sequences, to iteratively outpaint local regions in each step. It introduces two modules: the Visual-textual Consistency Refiner, which enhances input utilization during generation, and the Global-local Mamba Adapter, which ensures global coherence across the generated panorama. In contrast to methods leveraging pre-trained LDMs (such as AOG-Net and OPa-Ma), the recent PAR [25] adopts a text-conditioned masked autoregressive model [78] for 360-degree panoramic image outpainting. Guided by text descriptions, PAR incorporates its designed dual-space circular padding and a translation consistency loss to improve output quality.
IV-C Comparisons
@@
To systematically evaluate the strengths and weaknesses of representative methods of different kinds, we conduct both qualitative and quantitative benchmarks on Text-Only Generation (Sec. IV-A) and Text-Driven NFoV Outpainting (Sec. IV-B) methods with publicly available inference codes and model checkpoints. For Text-Only Generation, we compare Text2Light [56], Diffusion360 [58], StitchDiffusion [57] and PanFusion [59]. For Text-Driven NFoV Outpainting, we compare PanoDiff [42] and Diffusion360.
Figure 7: Visual comparison of text-driven NFoV outpainting approaches. The rest caption is the same as that for Fig. 6.
TABLE IV: Quantitative Comparison of Representative Text-Driven 360-Degree Panorama Generation. The first block of rows are for methods in the paradigm of Text-Only Generation, while the second block of rows are for Text-Driven NFoV Outpainting. We use metrics outlined in Sec. III-C for comprehensive evaluation. The inference time and GPU memory, required by each method to generate a 1024$\times$512 360-degree panorama, are also reported. The best and second-best results are highlighted for the two paradigms, respectively.
Method FID $\downarrow$ KID (×$\times 10^{-2}$) $\downarrow$ IS $\uparrow$ CS $\uparrow$ FAED $\downarrow$ OmniFID $\downarrow$ DS $\downarrow$ Inference (s) $\downarrow$ GPU Memory (GB) $\downarrow$
Text2Light 72.63 1.54 5.35 19.20 18.10 99.81 5.38 33 12.5
Diffusion360 70.32 2.00 5.29 18.74 12.43 92.23 0.94 3 3.5
StitchDiffusion 76.69 2.04 7.36 19.20 15.58 108.63 1.07 28 3.6
PanFusion 61.23 1.07 6.16 18.96 13.16 92.22 0.85 30 26.3
PanoDiff 65.94 2.44 4.72 19.02 10.24 122.30 1.10 48 36.0
Diffusion360 64.19 2.05 4.53 17.92 5.50 101.39 0.72 4 3.7
IV-C1 Qualitative Comparison
Fig. 6 presents a visual comparison of the four text-only generation methods. To facilitate the assessment of visual and semantic continuities at the image borders, we generate ”swapped” versions by exchanging the left and right halves of each generated image. Red dashed boxes are employed in these swapped images to highlight the boundary regions. Among the evaluated methods, Text2Light exhibits obvious seams in the highlighted areas of its swapped images, indicating a failure to maintain visual and semantic continuities at the borders. While Diffusion360 effectively mitigates such observable seams and preserves visual continuity, it occasionally demonstrates local semantic discontinuity at the borders, where content across image boundaries lacks consistent logical or meaningful coherence. In contrast, both StitchDiffusion and PanFusion achieve strong visual and semantic continuities at the borders of their synthesized images.
Fig. 7 displays the visual comparison for the two text-driven NFoV outpainting approaches. Similar to the text-only comparison, swapped image versions were created to scrutinize the visual and semantic continuities at the borders. Both PanoDiff and Diffusion360 demonstrate commendable visual and semantic continuities at the borders of the produced images.
IV-C2 Quantitative Comparison
To quantitatively assess the performance of these representative methods, we conduct a comprehensive comparison using metrics outlined in Sec. III-C. To ensure an unbiased evaluation of the generalizability of the methods, we employ an out-of-domain dataset, ODI-SR [1], on which none of the models have been explicitly trained. For generating text descriptions, we utilize BLIP2 [18] to create textual captions for the 360-degree panoramas included in the ODI-SR dataset. These generated text prompts serve as inputs for both Text-Only Generation and Text-Driven NFoV Outpainting tasks. To simulate NFoV images, we first project the equirectangular 360-degree panoramas from the ODI-SR dataset into a cubemap format and then extract the front face of each cubemap. The original 360-degree panoramas from ODI-SR are designated as real images and used as ground truth for the computation of evaluation metrics.
The quantitative results obtained from this comparative evaluation, across the seven evaluation metrics, are presented in Tab. IV. To provide insights into the computational efficiency of each method, we also report the inference time and GPU memory required to generate a 1024$\times$512 360-degree panorama on a consistent machine equipped with an RTX A6000 GPU. These quantitative findings effectively delineate the strengths and weaknesses of the evaluated approaches, offering valuable guidance for future research in this domain.
V Emerging 3D Applications
Recent advances in text-driven 360-degree panorama generation [57, 58, 59] have catalyzed innovative methods for reconstructing 360-degree 3D scenes from textual descriptions. 360-degree panoramic images inherently capture both global contexts and geometric constraints of a scene, making them an essential intermediate representation for 3D scene generation. Consequently, recent text-driven 360-degree 3D scene generation methods [31, 30, 33, 32, 27] utilize 360-degree panorama generation to bridge the gap between text prompts and 360-degree 3D scene reconstruction.